OpenStudy (anonymous):

Solve. 2x2 – 12x + 20 = 0 3 ± i 3 ± 2i 1 ± 2i 2 ± 3i

OpenStudy (anonymous):

is it the last one?

OpenStudy (anonymous):

Use the quadratic formula.

OpenStudy (anonymous):

use the quadratic formula....\(\huge x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \) where a, b, c are the coefficients of \(\large ax^2+bx+c=0 \)

OpenStudy (anonymous):

yea i know that but that little i thingy confuses me

OpenStudy (anonymous):

first simplify by 2 to get\[ 2x^2 – 12x + 20 = 0\\ 2( x^2 -6 x + 10) =0\\ x^2 -6 x + 10 =0\\ \] then apply the quadratic formula as @dpalnc suggested

OpenStudy (agent47):

i=sqrt(-1)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

a=1, b=-6, c=10\\ \[\Delta = b^2 -4 ac = 36 -40= -4 \\ \sqrt{\Delta} = \pm 2 i \]

OpenStudy (anonymous):

the first choice?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[ \frac {6\pm 2i}{2}=3\pm i \]

OpenStudy (anonymous):

yea i got that.. ty

OpenStudy (anonymous):

yw